If instantaneous current is I=2cos(omega t+phi) ampere, then rms value of current is?

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

The peak value of the current isI_0=2A
Hence rms value of current=\dfrac{I_0}{\sqrt{2}}=\sqrt{2}A